I need steps for simplifying 6(sqrt7/sqrt4)+3(sqrt28)-10(sqrt1/sqrt7).
Thanks so much!

Found 2 solutions by dabanfield, checkley77:
Answer by dabanfield(803)

I need steps for simplifying 6(sqrt7/sqrt4)+3(sqrt28)-10(sqrt1/sqrt7).
6(sqrt7/sqrt4)+3(sqrt28)-10(sqrt1/sqrt7) =
6*(sqrt7)/2) + 3*(sqrt(4)*sqrt(7)) - 10*(1/sqrt(7)) =
(6/2)*sqrt(7) + 3*2*sqrt(7) - 10/sqrt(7) =
3*sqrt(7) + 6*sqrt(7) - 10/sqrt(7) =
9*sqrt(7) - 10/sqrt(7)
To remove the "sqrt" from the denominator of the second term we can multiply both its numerator and denominator by sqrt(7):
9*sqrt(7) - (10*sqrt(7))/((sqrt(7)*sqrt(7)) =
9*sqrt(7) - 10*sqrt(7)/7 =
63/7*sqrt(7) - 10/7*sqrt(7) =
(53/7)*sqrt(7)

Answer by checkley77(12844)

6(sqrt7/sqrt4)+3(sqrt28)-10(sqrt1/sqrt7) [factor sqrt28=sqrt4*7=2sqrt7, sqrt4=2 &sqrt1=1.
6(sqrt7)/2+3*2(sqrt7)-10/(sqrt7)
3(sqrt7)+6(sqrt7)-10/(sqrt7) find a common denominator (sqrt7).
[3(sqrt7)(sqrt7)+6(sqrt7)(sqrt7)-10]/(sqrt7)
[3*7+6*7-10]/(sqrt7)
[21+42-10]/(sqrt7)
[63-10]/(sqrt7)
53/(sqrt7) ans.